# Read, Watch and Scream! Sound Generation from Text and Video

### v1

@@ Abstract
Multimodal generative models have shown impressive advances with the help of powerful diffusion models. Despite the progress, generating sound solely from text poses challenges in ensuring comprehensive scene depiction and temporal alignment. Meanwhile, video-to-sound generation limits the flexibility to prioritize sound synthesis for specific objects within the scene. To tackle these challenges, we propose a novel video-and-text-to-sound generation method, called ReWaS, where video serves as a conditional control for a text-to-audio generation model. Our method estimates the structural information of audio (namely, energy) from the video while receiving key content cues from a user prompt. We employ a well-performing text-to-sound model to consolidate the video control, which is much more efficient for training multimodal diffusion models with massive triplet-paired (audio-video-text) data. In addition, by separating the generative components of audio, it becomes a more flexible system that allows users to freely adjust the energy, surrounding environment, and primary sound source according to their preferences. Experimental results demonstrate that our method shows superiority in terms of quality, controllability, and training efficiency. Our demo is available at https://naver-ai.github.io/rewas.
@@ 1 Introduction
Recent generative models have developed dramatically, making content creation easier for people, including images, videos, and audio, based on text prompts. Especially, text-to-video generation models such as Make-a-Video [35] and Sora [1] show the impressive emergence of generative models in the video domain, showing remarkable utility in film and advertising. While we are fully immersed in the video content by watching and listening, unfortunately, these generated videos are silent. Generating the sound of a video is a challenging task requiring both a contextual and temporal understanding of the video. Figure 1 shows an actual example of when both text and video controls are required for generating realistic audio for the given video. In the video, a dog is growling while holding a toy in his mouth. A human can imagine the sound of the video; the dog growls lowly, and the growling sounds like the dog is biting something. When the person grips and pulls the toy, the dog will treat the human by growling louder. Finally, when the dog shakes his head, the growling will become louder. If the generated audio does not understand the visual information, it will be a random growling sound, and it will not sound like the dog bites something. If the audio is not controlled by the text, the generated audio might be only related to the dog, e.g., a barking sound.
@@
Table 1 shows the recent attempts to generate an audio sample from the given video or the given text context. As shown in the table, there are two major directions to generate an audio sample from the given video directly. For example, there have been studies of a sound effect (SFX) generation with short moments for video editing tasks [4, 7, 43], known as Foley. These methods are restricted to the pre-defined sound effect classes and can only control discrete information, such as onset. As another attempt, video-to-audio (V2A) generation methods have been proposed recently [27, 44, 18, 34]. They hardly generate sounds from multiple objects together but likely generate only simple sounds. Furthermore, both SFX and V2A methods cannot take text controls, which provide more rich control by users. Figure 1 shows the example when there is no text control; a V2A method just generates audio of “barking” rather than “growling” by focusing on the dog in the video. Our work aims to generate general sounds conditioned on both video and text for user control.
Figure 1: An example of audio generation requiring both text and video control. The text instruction “dog growling” is used for the text control. The video-to-audio (V2A) [34] or text-to-audio (T2A) [25] generation methods cannot understand the detailed semantics from texts (the dog is growling, not barking) or video (the dog is biting something, and the alignment), respectively.
@@
As another line of research, text-to-audio (T2A) generation has been actively studied [17, 25, 26, 10, 16]. Despite their diverse and high-quality audio generation quality, they lack a temporal understanding of video-only information. For example, like the example in Figure 1, text-only condition can make irrelevant audio to the video, namely, it might not be aligned with the given video, and it might not consider information only implied by visual information (e.g., the dog holds something in the mouth). To tackle the problem, we may need more controllability to the T2A model, such as AudioLDM [25]. Recently, a few studies [41, 12] tried to control the pre-trained AudioLDM more precisely based on ControlNet [46]. Although they can control the pitch, temporal order, energy, or rhythm of the generated audio, their generation process needs expensive timestamp-wise annotations for each control feature. Namely, the existing T2A works cannot reflect the visual information hidden from the given text or pre-defined control features, and requires expensive timestamp-wise control features.
In this work, we propose a novel video-and-text-to-sound generation approach, named Read, Watch and Scream (ReWaS), by integrating video as a conditional control for a well-established text-to-audio model. Our method is based on a state-of-the-art text-to-audio generation method, AudioLDM [25]. While a text prompt specifies the subject, we additionally employ a control feature extracted from the video. More specifically, our method presents an energy adapter on AudioLDM motivated from ControlNet [46], an efficient structure control method for text-to-image generation. Since a video feature does not directly imply the structure of the audio, we estimate the temporal energy information, a basic audio structural information, from the video.
The energy operates as a time-varying control to complement the sound according to the dynamics of the given video. As shown in Figure 1, ReWaS successfully understands complex information from both text and video. Here, we define energy as the mean of frequency in each audio frame, which is related to both visual dynamics and semantics [20, 12]. It is relatively simple to estimate from a video rather than complex acoustic features (e.g., MFCC, mel-spectrograms). Therefore, our energy control facilitates connecting video for T2A model. Lastly, our method offers a more flexible and effective solution thanks to the efficient training [46].
We compare our method and other state-of-the-art video-to-audio generation models [7, 43, 27, 18, 34] on two video-audio aligned datasets, VGGSound [2] and GreatestHits [28]. In the experiments, ReWaS shows a superior audio generation performance quantitatively and qualitatively. For example, our method shows the best fidelity score (FID) and structure prediction (energy MAE) on both datasets. Furthermore, ReWaS outperforms V2A methods in human evaluation for three categories (audio quality, relevance to the video, and temporal alignment between audio and video) with a significant gap (almost +1 point for every category in 5-scale MOS). Also, as shown in the qualitative study, our method successfully deals with the temporal alignment of the visual information. For example, our method can capture the “short transition” of the skateboarding video when the boarder jumps into the air, and no skateboarding sound appears in the video.
2 Related work
Table 1: Comparison of audio generation methods. We consider four factors: Can it make a general sound? Can it take text or visual control? and is its training efficient?
Method General sound? Text control? Visual control? Efficient training?
Sound effect (SFX) generation [4, 7, 43]
Video-to-audio (V2A) [27, 44, 18, 34]
Text-to-audio (T2A) [17, 25, 26, 10, 16]
T2A + Control [41, 12]
ReWaS (ours)
@@ 2.1 Text-to-audio generation
Early work for conditional audio generation was built upon generative adversarial networks [24, 6], normalizing flows [21], and variational autoencoders [39]. Recently, several studies based on diffusion models have shown promising progress on a broad range of acoustic domains. DiffSound [45] employs a diffusion-based token decoder for the first time to transfer text features into mel-spectrogram tokens. Make-An-Audio [17], AudioLDM [25], AudioLDM2 [26], Tango [10] and Make-An-Audio2 [16] are well-founded in latent diffusion model (LDM) [32], demonstrating high-quality results with large scale training. A series of LDM predicts mel-spectrograms using a VQ-VAE decoder, and a pretrained vocoder generates raw waveforms from the generated mel-spectrograms. While these methods successfully generate high-quality audio samples for the given text prompt, they are only designed for taking text conditions, unable to understand visual semantics.
Meanwhile, there have been a few attempts based on ControlNet [46], an efficient training method for structure control for text-to-image generation. ControlNet utilizes hints (e.g., Canny edge maps, scribbles, human pose, depth maps) to provide structural composition to the generated images. Inspired by this, MusicControlNet [41] showed control over melody, dynamics, and rhythm, while Guo et al. [12] built a FusionNet between each layer of the U-Net, enabling the fusion of control embeddings for temporal order, pitch, and energy controls. They have demonstrated that incorporating control signals into the pretrained audio generative models provides more explicit and fine-grained control over the generated audio, leading to performance improvement and adherence to the desired characteristics. However, designing these time-varying controls still requires skilled labor for users. To address this challenge, we generate energy control through a given video, which is a practical function for creating SFX, post-production for filmmaking, and utilizing AI-generated silent videos.
@@ 2.2 Video-to-audio generation
Existing video-to-audio (V2A) generation methods have focused on achieving two main characteristics: (i) audiovisual relevance and (ii) temporal synchronization. The first stream aims to represent general sound by leveraging datasets such as VGGSound [2] and AudioSet [9]. Given a set of video features, SpecVQGAN [18] learns a transformer to sample quantized representations (i.e., codebook) based on visual features to decode spectrogram. Im2wav [34] utilizes rich semantic representations obtained from a pre-trained CLIP [30] as sequential visual conditioning for an audio language model, and applies classifier-free guidance [15] to steer the generation process. Recently, diffusion-based models have shown the stunning ability to generate high-quality audio [27, 44]. DiffFoley [27] is a diffusion-based video-to-audio generation model that learns temporal and semantic alignment features through contrastive learning. Although it achieves better alignment between visual and audio inputs with prior training, it necessitates tremendous training data, such as the utilization of both VGGSound and AudioSet for alignment training. Seeing-and-hearing [44] is another diffusion-based model that optimizes the generation process using ImageBind [11] which learns joint embedding space for six modalities (image, text, audio, depth, thermal, and IMU). However, they often struggle to generate temporally aligned sounds at short times in the video (e.g., dog barking, people laughing).
On the other hand, other research works [4, 43] have focused on creating simplistic SFX (e.g., stick hits) using datasets like CountixAV [47] and GreatestHits [28], which provide fewer classes but more precisely temporal aligned data. CondFoleyGen [7] trains a transformer model to autoregressively predict a sequence of audio codes for a spectrogram VQGAN, conditioned on the given audiovisual example. Syncfusion [4] predicts a discrete onset label that denotes the beginning of a sound for repetitive actions. Recent SonicVisionLM [43] employs a large language model (LLM) to utilize text as an intermediate product that facilitates user interaction for personalized sound generation. They freeze Tango [10] and train ControlNet with timestamp estimated by a video for 23 SFX categories exclusively, where a sound event timestamp detection module is trained on a dataset comprising videos and sound timing data. While these approaches are limited to generating only pre-defined training SFX categories, our method generates sounds for various categories from the visual context.
@@ 3.1 Text-to-audio latent diffusion model
In this paper, we specifically utilize AudioLDM [25] which generates a latent of mel-spectrogram $z$ computed by VAE [22]. The diffusion model $\epsilon_{\theta}$ of AudioLDM is trained to predict the noise added to a given data by minimizing the objective function, $\mathcal{L}_{\text{diff}}=\mathbb{E}_{z_{0},\epsilon,t}\left\|\epsilon-\epsilon_{\theta}(z_{t},t,\mathbf{E}_{a})\right\|^{2}_{2}$, where $\epsilon$ represents the noise added at time $t$, $z_{t}$ is noisy latent induced via the forward process and $\mathbf{E}_{a}$ denotes the embedding of the audio $x$ obtained from the CLAP audio encoder $f_{\text{audio}}(\cdot)$ [42]. Here, the model is conditioned by $\mathbf{E}_{a}$ using classifier free guidance [15].
@@
In the sampling process, the generation starts from a noise $z_{T}$ sampled from $\mathcal{N}(0,I)$ and the text embedding $\mathbf{E}_{y}$ from the CLAP text encoder $f_{\text{text}}(\cdot)$. The reverse process conditioned on $\mathbf{E}_{y}$ gradually generates the audio prior $z_{0}$ using the modified noise estimation $\hat{\epsilon}_{\theta}(z_{t},t,\mathbf{E}_{y})=(1+w)\epsilon_{\theta}(z_{t},t,\mathbf{E}_{y})-w\epsilon_{\theta}(z_{t},t)$, where $w\in[0,1]$ is a guidance weight to balance the trade-off of the audio condition $\mathbf{E}_{a}$. The VAE decoder decodes the sampled latent $z$ to predict a mel-spectrogram. Finally, the decoded mel-spectrogram is converted to a raw audio sample using the HiFi-GAN vocoder [23].
@@
Although AudioLDM enables text-conditional audio generation, it still lacks of understanding of visual contents and their temporal information. In this paper, we add a visual control to the pre-trained AudioLDM. Instead of directly using a visual feature to control, we extract more essential information from the given video, which will be discussed in Section 3.2.
3.2 Rethinking video-to-sound generation with temporal alignment
Figure 2: Limitation of timestamp annotations.
Figure 3: Energy can imply multiple semantics.

### v2

@@ Abstract
Despite the impressive progress of multimodal generative models, video-to-audio generation still suffers from limited performance and limits the flexibility to prioritize sound synthesis for specific objects within the scene. Conversely, text-to-audio generation methods generate high-quality audio but pose challenges in ensuring comprehensive scene depiction and time-varying control. To tackle these challenges, we propose a novel video-and-text-to-audio generation method, called ReWaS, where video serves as a conditional control for a text-to-audio generation model. Especially, our method estimates the structural information of sound (namely, energy) from the video while receiving key content cues from a user prompt. We employ a well-performing text-to-audio model to consolidate the video control, which is much more efficient for training multimodal diffusion models with massive triplet-paired (audio-video-text) data. In addition, by separating the generative components of audio, it becomes a more flexible system that allows users to freely adjust the energy, surrounding environment, and primary sound source according to their preferences. Experimental results demonstrate that our method shows superiority in terms of quality, controllability, and training efficiency. Code and demo are available at https://naver-ai.github.io/rewas
Figure 1: An example of audio generation requiring both text and video control. The text instruction “dog growling” is used for the text control. The video-to-audio (V2A) (Sheffer and Adi 2023) or text-to-audio (T2A) (Liu et al. 2023) generation methods cannot understand the detailed semantics from texts (the dog is growling, not barking) or video (the dog is biting something, and the alignment), respectively.
11footnotetext: Work done during an internship at NAVER AI Lab.22footnotetext: Corresponding author: lee.j@navercorp.com.
@@ 1 Introduction
Generative models have developed dramatically, making content creation easier for people. Especially, text-to-video generation models such as Make-a-Video (Singer et al. 2022) and Sora (Brooks et al. 2024) show the impressive emergence of generative models in the video domain, showing remarkable utility in film and advertising. While we are fully immersed in the video content by watching and listening, unfortunately, these generated videos are silent. Generating the sound aligned to a video is a challenging task requiring both a contextual and temporal understanding of the video. Figure 1 shows an example of when text and video controls are required to generate a sound that precisely matches the given scene. Here, the dog is growling while holding a toy in his mouth. A human can imagine the sound of the video; the dog growls lowly, and the growling sounds like the dog is biting something. When the person grips and pulls the toy, the dog will treat the human by growling louder. Finally, when the dog shakes his head, the growling will become louder. If a generative model does not understand the visual information, it will be a random growling sound, not like the dog biting something. If audio is not controlled by text, the generated audio might be only related to the dog, e.g., a barking sound.
@@
Table 1 shows the recent attempts to generate an audio sample from the given video or text. There are two major directions to generate an audio sample from the given video directly. First, there have been studies of a sound effect (SFX) generation with short moments for video editing tasks (Comunità et al. 2024; Du et al. 2023), known as Foley. They are restricted to the pre-defined sound effect classes and can only control discrete information, such as onset. As another attempt, video-to-audio (V2A) generation methods have been proposed (Luo et al. 2024; Xing et al. 2024; Iashin and Rahtu 2021; Sheffer and Adi 2023). However, they still struggle to generate open-domain sounds from multiple objects together. Furthermore, both SFX and V2A methods cannot take text controls, more rich user control. Figure 1 shows the example when there is no text control; a V2A method just generates audio of “barking” rather than “growling” by focusing on the dog in the video.
@@
As another line of research, text-to-audio (T2A) generation has been actively studied (Huang et al. 2023a, b; Liu et al. 2023, 2024; Ghosal et al. 2023). Despite their diverse and high-quality audio generation quality, they lack a temporal understanding of video-only information. Like the example in Figure 1, the text-only condition can make irrelevant audio to the video (e.g., when the dog shakes heads). To tackle the problem, we may need more controllability to the T2A model, such as AudioLDM (Liu et al. 2023). Recently, a few studies (Wu et al. 2023; Guo et al. 2024a; Chung, Lee, and Nam 2024) tried to control the pre-trained AudioLDM more precisely based on ControlNet (Zhang, Rao, and Agrawala 2023). Although they can control the pitch, temporal order, energy, or rhythm of the generated audio, their generation process needs expensive timestamp-wise annotations for each control feature.
More recently, parallel to our study, SonicVisionLM (Xie et al. 2024) and Seeing&Hearing (Xing et al. 2024) incorporate text information, providing users the freedom to generate specific sounds. Although these methods can control audio generation with both vision and language, they still suffer from either limited discrete control (e.g., onset) (Xie et al. 2024), or lacking timestamp-wise control (Xing et al. 2024). Moreover, they require a video-to-text converting process, such as video captioning or feature mapping, for use with the T2A model. This text conversion weakens temporal alignment, leading to the loss of fine-grained temporal details.
Method General Text Visual W/o V2T Efficient
sound? control? control? mapping? training?
Sound effect generation (Comunità et al. 2024; Du et al. 2023)
Video-to-audio (Luo et al. 2024; Iashin and Rahtu 2021; Sheffer and Adi 2023)
Text-to-audio (Huang et al. 2023a, b; Liu et al. 2023, 2024; Ghosal et al. 2023)
Text-to-audio & Control (Wu et al. 2023; Guo et al. 2024a; Chung, Lee, and Nam 2024)
Video-to-text & Text-to-audio (Xie et al. 2024; Xing et al. 2024)
Video-and-text-to-audio (ReWaS)
@@
  • $\star$
    Taking limited timestamp-wise visual control (e.g., requiring the full timestamp-wise onset annotations, or only able to take a few frames)
Table 1: Comparison of audio generation methods: Can it make a general sound? Can it take text or visual control? Does it need video-to-text (V2T) mapping? and the training efficiency.
In this work, we propose a novel video-and-text-to-audio generation approach, named Read, Watch and Scream (ReWaS), by integrating video as a conditional control for a well-established T2A model. While a text prompt specifies the subject, we additionally employ a control feature extracted from the video. More specifically, our method presents an energy adapter on AudioLDM motivated from ControlNet (Zhang, Rao, and Agrawala 2023), an efficient structure control method for text-to-image generation. Since a video feature does not directly imply the structure of the audio, we estimate the temporal energy information, a basic audio structural information, from the video.
The energy operates as a time-varying control to complement the sound according to the dynamics of the given video. As shown in Figure 1, ReWaS successfully understands complex information from both text and video. Here, we define energy as the mean of frequency in each audio frame, which is related to visual dynamics and semantics (Jeong et al. 2023; Guo et al. 2024a). It is relatively simple to estimate from a video rather than complex acoustic features (e.g., mel-spectrograms). Therefore, our energy control facilitates connecting video for T2A model, reflecting strong alignment between audio and video.
We compare our method and other state-of-the-art video-to-audio generation models (Du et al. 2023; Xing et al. 2024; Luo et al. 2024; Iashin and Rahtu 2021; Sheffer and Adi 2023) on two video-audio aligned datasets, VGGSound (Chen et al. 2020) and GreatestHits (Owens et al. 2016). In the experiments, ReWaS outperforms V2A methods in human evaluation for three categories (audio quality, relevance to the video, and temporal alignment between audio and video) with a significant gap (almost +1 point for every category in 5-scale MOS). Also, ReWaS shows a superior audio generation performance quantitatively and qualitatively. Our method shows the best fidelity score (FD), structure prediction (energy MAE), and AV-alignment score on VGGSound. Moreover, we achieve the best AP and energy MAE on Greatest Hits without the use of reference audio samples like CondFoleyGen (Du et al. 2023). As shown in the qualitative study, ReWaS can capture the challenging “short transition” of the video when the skateboarder jumps into the air, and no skateboarding sound appears in the video. It is also possible to generate the sound of a video generated by a general text prompt.
2 Related Work
@@ 2.1 Text-to-audio generation
Early work for audio generation was built upon GANs (Kreuk et al. 2023; Dong et al. 2018), normalizing flows (Kim et al. 2020), and VAEs (Van Den Oord, Vinyals et al. 2017). Recently, several studies using diffusion models have shown promising progress on a broad range of acoustic domains. DiffSound (Yang et al. 2023) employs a diffusion-based token decoder for the first time to transfer text features into mel-spectrogram tokens. Make-An-Audio (Huang et al. 2023b), AudioLDM (Liu et al. 2023), AudioLDM2 (Liu et al. 2024), Tango (Ghosal et al. 2023) and Make-An-Audio2 (Huang et al. 2023a) are well-founded in latent diffusion model (LDM) (Rombach et al. 2022), demonstrating high-quality results with large scale training. A series of LDM predicts mel-spectrograms using a VQ-VAE decoder, and a pretrained vocoder generates raw waveforms from the generated mel-spectrograms. While these methods successfully generate high-quality audio samples for the given text prompt, they are only designed for taking text conditions, unable to understand visual semantics.
Meanwhile, there have been a few attempts based on ControlNet (Zhang, Rao, and Agrawala 2023), an efficient training method for structure control for text-to-image generation. ControlNet utilizes hints (e.g., Canny edge maps, scribbles, depth maps) to provide a structural composition to the generated images. Inspired by this, text-to-audio methods have incorporated ControlNet to accomplish controllable music (Wu et al. 2023) and audio effect (Guo et al. 2024a; Chung, Lee, and Nam 2024) generation. They have provided more explicit and fine-grained control over the generated audio, leading to performance improvement and adherence to the desired characteristics.
However, designing these time-varying controls still requires costly labor for users. To address this challenge, we predict energy control through a given video, which is a practical function for creating SFX, post-production for filmmaking, and utilizing AI-generated silent videos.
@@ 2.2 Video-to-audio generation
Existing video-to-audio (V2A) generation methods have focused on achieving two main characteristics: (i) audiovisual relevance and (ii) temporal synchronization. The first stream aims to represent general sound by leveraging datasets such as VGGSound (Chen et al. 2020) and AudioSet (Gemmeke et al. 2017). Given a set of video features, SpecVQGAN (Iashin and Rahtu 2021) learns a transformer to sample quantized representations (i.e., codebook) based on visual features to decode spectrogram. Im2wav (Sheffer and Adi 2023) utilizes rich semantic representations obtained from a pre-trained CLIP (Radford et al. 2021) as sequential visual conditioning for an audio language model, and applies CFG (Ho and Salimans 2022) to steer the generation process. Recently, diffusion-based models have shown the stunning ability to generate high-quality audio (Luo et al. 2024; Xing et al. 2024). DiffFoley (Luo et al. 2024) improves audiovisual relevance by learning temporal and semantic alignment through contrastive learning. However, it necessitates tremendous training data, such as the utilization of both VGGSound and AudioSet for alignment training. Seeing&hearing (Xing et al. 2024) is another diffusion-based model that optimizes the text-to-audio diffusion model, AudioLDM (Liu et al. 2023) by using ImageBind (Girdhar et al. 2023) which learns joint embedding space for six modalities (image, text, audio, depth, thermal, and IMU). However, ImageBind Video Encoder takes only two frames for each video sampled from 2 second, which results in lacking timestamp-wise contol. Therefore, they often struggle to generate temporally aligned sounds at short times in the video (e.g., dog barking, people laughing).
On the other hand, other research works (Comunità et al. 2024; Xie et al. 2024) have focused on creating simplistic SFX (e.g., stick hits) using datasets like CountixAV (Zhang, Shao, and Snoek 2021) and GreatestHits (Owens et al. 2016), which provide fewer classes but more precisely temporal aligned data. CondFoleyGen (Du et al. 2023) trains a Transformer to autoregressively predict a sequence of audio codes for a spectrogram VQGAN, conditioned on the given audiovisual example. Syncfusion (Comunità et al. 2024) predicts a discrete onset label that denotes the beginning of a sound for repetitive actions. Recent SonicVisionLM (Xie et al. 2024) employs a large language model to utilize text as an intermediate product that facilitates user interaction for personalized sound generation. They freeze Tango (Ghosal et al. 2023) and train ControlNet with timestamp estimated by a video for 23 SFX categories exclusively, where the video is converted to sound event timestamp and text. Although they have shown promising results in SFX generation, their timestamp detection module is limited to a single visual object, and they cannot implicit detailed temporal cues in visual content because they use videos to convert them into text. our method generates sounds for various categories from the visual context at the same time.
@@ 3.1 Text-to-audio latent diffusion model
In this paper, we specifically utilize AudioLDM (Liu et al. 2023) which generates a latent of mel-spectrogram $z$ computed by VAE (Kingma and Welling 2014). The diffusion model $\epsilon_{\theta}$ of AudioLDM is trained to predict the noise added to a given data by minimizing the objective function, $\mathcal{L}_{\text{diff}}=\mathbb{E}_{z_{0},\epsilon,t}\left\|\epsilon-\epsilon_{\theta}(z_{t},t,\mathbf{E}_{a})\right\|^{2}_{2}$, where $\epsilon$ represents the noise added at time $t$, $z_{t}$ is noisy latent induced via the forward process and $\mathbf{E}_{a}$ denotes the embedding of the audio $x$ obtained from the CLAP audio encoder $f_{\text{audio}}(\cdot)$ (Wu* et al. 2023). Here, the model is conditioned by $\mathbf{E}_{a}$ using classifier free guidance (CFG) (Ho and Salimans 2022).
@@
In the sampling process, the generation starts from a noise $z_{T}$ sampled from $\mathcal{N}(0,I)$ and the text embedding $\mathbf{E}_{y}$ from the CLAP text encoder $f_{\text{text}}(\cdot)$. The reverse process conditioned on $\mathbf{E}_{y}$ generates the audio prior $z_{0}$ using the modified noise estimation $\hat{\epsilon}_{\theta}(z_{t},t,\mathbf{E}_{y})=(1+w)\epsilon_{\theta}(z_{t},t,\mathbf{E}_{y})-w\epsilon_{\theta}(z_{t},t)$, where $w$ is a guidance weight to balance the audio condition $\mathbf{E}_{a}$. The VAE decoder decodes the sampled latent $z$ to predict a mel-spectrogram. Finally, the decoded mel-spectrogram is converted to a raw audio sample using the HiFi-GAN vocoder (Kong, Kim, and Bae 2020).
@@
Although AudioLDM enables text-conditional audio generation, it still lacks of understanding of visual contents and their temporal information. This study adds a visual control to the pre-trained AudioLDM. Instead of directly using a visual feature to control, we extract more essential information from the given video, which will be discussed in Section 3.2.
3.2 Video-to-audio with temporal alignment
Figure 2: Discrete timestamp annotations vs. Continuous energy.
Figure 3: Energy can improve temporal alignment.